OpenStudy (anonymous):

If f(x) = 5x^2 − 4 and g(x) = x − 2, what is (f o g)(-3)?

OpenStudy (anonymous):

igot f(x)=41 & g(x)=-5, but that isn't a choice for my answer?

OpenStudy (mathstudent55):

(fog)(x) = f(g(x)) First evaluate g(-3) in g(x) = x - 2 Then the number you get from that substitute for x in f(x) = 5x^2 - 4

OpenStudy (anonymous):

k, i got 121

OpenStudy (anonymous):

that is an answer

OpenStudy (mathstudent55):

g(-3) = -3 - 2 = -5 f(-5) = 5(-5)^2 - 4 = 5(25) - 4 = 121

OpenStudy (anonymous):

Thanks for your help again....i got to sign off now , but I greatly appreciate your help and instruction